A new vaccination is being used in a laboratory experiment to investigate whether it is effective. There are 203 subjects in the study. Is there sufficient evidence to determine if vaccination and disease status are related?

Vaccination Status   Diseased   Not Diseased   Total
Vaccinated   42   37   79
Not Vaccinated   58   66   124
Total   100   103   203

Step 8 of 8: State the conclusion of the hypothesis test at the 0.05 level of significance.
Answer #1

Answer:

There is not sufficient evidence to conclude that vaccination and disease status are related.

Solution:

For this Chi-square test, the null and alternative hypotheses are given as below:

Null hypothesis: H0: vaccination and disease status are not related.

Alternative hypothesis: Ha: vaccination and disease status are related.

Test statistic formula is given as below:

Chi square = ∑[(O – E)^2/E]

Where, O is observed frequencies and E is expected frequencies.

E = row total * column total / Grand total

Calculation tables for test statistic are given as below:

Observed Frequencies

Column variable

Row variable

Diseased

Not Diseased

Total

Vaccinated

42

37

79

Not Vaccinated

58

66

124

Total

100

103

203

Expected Frequencies

Column variable

Row variable

Diseased

Not Diseased

Total

Vaccinated

38.91626

40.08374

79

Not Vaccinated

61.08374

62.91626

124

Total

100

103

203

Calculations

(O - E)

3.083744

-3.08374

-3.08374

3.083744

(O - E)^2/E

0.244357

0.23724

0.155679

0.151145

Test Statistic = Chi square = ∑[(O – E)^2/E] = 0.788422

χ2 test statistic = 0.788

Step 7 of 8:

Make the decision to reject or fail to reject the null hypothesis at the 0.05 level of significance.

α = 0.05

We are given

Number of rows = r = 2

Number of columns = c = 2

Degrees of freedom = df = (r – 1)*(c – 1) = 1*1 = 1

Critical value = 3.841459

(by using Chi square table or excel)

P-value = 0.374578

(by using Chi square table or excel)

P-value > α = 0.05

So, we fail to reject the null hypothesis

Answer:

Fail to reject the null hypothesis

Step 8 of 8: State the conclusion of the hypothesis test at the 0.05 level of significance.

We fail to reject the null hypothesis that vaccination and disease status are not related.

There is sufficient evidence to conclude that vaccination and disease status are not related.

There is not sufficient evidence to conclude that vaccination and disease status are related.

Answer:

There is not sufficient evidence to conclude that vaccination and disease status are related.